OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (anonymous):

@UditKulka

OpenStudy (anonymous):

@ganeshie8 @sarahg95

OpenStudy (anonymous):

@John_ES

OpenStudy (anonymous):

i think its 10

OpenStudy (john_es):

First, calculate the inverse function, \[y=2x+2\Rightarrow x'=2y'+1\Rightarrow y'=(x'-1)/2\]So\[f^{-1}(x)=(x-1)/2\] With x=4, \[f^{-1}(4)=3/2\]

OpenStudy (anonymous):

@John_ES thanks :)